Given the training data in Question 1 below| (on buying RRSP8), predict the class of the following new example using k-nearest-neighbor classification fork = 5: sector = oil industry, income = medium, self-employed = yes, credit-rating fair. For distance measure, use the following similarity measure: similarity(tupleAtupleB)-4-.(w"S(ab/4), where S(ab) is 1 if parameter a equals parameter b and o otherwise The parameters atand biare either Sector, income, self-employed, or credit-rating. The weights wiare all 1, except for income, which is 2. Please show your calculations or explain the steps. Note: the similarity measure is the opposite of the distance measure.
Sector of activity Self-Employed Credit-rating Person ID Class: Buys-RRSP Income farming medium fair 1 no no farming fair low 2 yes no oil industry medium fair no no oil industry low fair yes 4 yes oil industry medium excellent 5 yes yes banking medium excellent yes no oil industry high fair 7 no no oil industry high excellent no no banking high fair yes no farming low excellent 10 yes yes banking low excellent 11 yes yes farming medium fair 12 yes no high banking fair 13 yes yes farming medium excellent 14 no yes
Answer #1

Given K =5

And We have to predict for : sector = oil industry, income = medium, self-employed = yes, credit-rating fair

Also the similarity measure is given as : Similarity(tupleA, tupleB)= Wi (S(ai, b)) i=1

where S(a, b) =1 if parameter "a" equals parameter "b" and =0 otherwise The parameters a(i), b(i) are either Sector, income, self-employed, or credit-rating. The weights w(i) are all 1, except for income, which is 2

NOTE: For calculating similarity measure we just use the given formula and corresponding value of S and W, depending on the case if they match our training set and the query or not.

Person ID Sector of Activity Income Self-employed Credit Rating Similarity Measure Value
1 farming medium No Fair 0+2+0+1 = 3
2 farming low yes Fair 0+0+1+1 = 2
3 oil industry medium No Fair 1+2+0+1 = 4
4 oil industry low yes fair 1+0+1+1 = 3
5 oil industry medium yes excellent 1+2+1+0 = 4
6 banking medium no excellent 0+2+0+0 = 2
7 oil industry high no fair 1+0+0+1 = 2
8 oil industry high no excellent 1+0+0+0 = 1
9 banking high no fair 0+0+0+1 = 1
10 farming low yes excellent 0+0+1+0 = 1
11 banking low yes excellent 0+0+1+0 = 1
12 farming medium yes fair 0+2+1+1 = 1
13 banking high yes fair 0+0+1+1 = 2
14 farming medium no excellent 0+2+0+0 = 2

Person ID

Sector of Activity

Income

Self-

employed

Credit Rating

Similarity

Measure

Value

Rank

Highest

Similarity

Measure

Can we include it in (K=5)

Nearest Neighbour

Y=Category

Of Nearest

Neighbour

1

farming

medium

No

Fair

0+2+0+1 = 3

3

YES

No

2

farming

low

yes

Fair

0+0+1+1 = 2

5

YES

No

3

oil industry

medium

No

Fair

1+2+0+1 = 4

1

YES

No

4

oil industry

low

yes

fair

1+0+1+1 = 3

4

YES

Yes

5

oil industry

medium

yes

excellent

1+2+1+0 = 4

2

YES

Yes

6

banking

medium

no

excellent

0+2+0+0 = 2

6

NO

Yes

7

oil industry

high

no

fair

1+0+0+1 = 2

8

NO

No

8

oil industry

high

no

excellent

1+0+0+0 = 1

10

NO

No

9

banking

high

no

fair

0+0+0+1 = 1

11

NO

Yes

10

farming

low

yes

excellent

0+0+1+0 = 1

12

NO

Yes

11

banking

low

yes

excellent

0+0+1+0 = 1

13

NO

Yes

12

farming

medium

yes

fair

0+2+1+1 = 1

14

NO

No

13

banking

high

yes

fair

0+0+1+1 = 2

8

NO

Yes

14

farming

medium

no

excellent

0+2+0+0 = 2

9

NO

Yes

So we can see that we have 3 No and 2 yes so our final answer class label will be "No".

(Note: here we had to choose one among several choices for the fifth neighbour so we selected one randomly. and there will be a change of answer if we had chose another one. So in such case we just select a random one and make our decision of the class)